last monthly update: jun18
The performance of the various receivers
measured with the heiles calibration scans are listed by receiver
and epoch. The plots show the average of polA and polB. There are 3
sets of plots:
For receivers that cover a wide range of frequencies, the data will
be plotted with all frequencies combined, and then by frequency.
- Gain [K/Jy], Tsys[K], SEFD[Jy/Tsys], and average beam
- coma parameter, 1st sidelobe height, main beam
efficiency, and main beam + first sidelobe efficiency.
- Pointing errors [Asecs] vs az and za.
For the alfa receiver the data is
first plotted with all pixels together. The color is used to code
the pixel number. After the combined plot there are 7 more plots
showing each pixel by itself. For these plots the color is used to
distinguish the source used.
There are sections that show the cumulative system performance
and the system performance by month.
Links to sections:
system performance by month.
of scans by rcvr for last 12 months (.ps) (.pdf)
flux of the calibration sources
Miscellaneous multi frequency
variations before/after hurricane maria.
This is a short list of changes that could affect the system
performance. A more complete list can be found at system
modification history: (top)
- 11dec17: added 05-07 model data offsets correctly for
cb,xb,430. (mainly affected 430).
- 09dec17: updated model19 offsets for 430,cband,xband including
- (but did it wrong.. 05-07 data had model19 offsets. mainly
- 04-07dec17. recheck other rcvr offsets with model19 installed.
- 01dec17: installed model19
- 21-25nov17: verify model19. use same offset as rcv had with
sbn in model18
- 19nov17: svn with model19, but platform way out of position.
- 01-17nov17: td tracking. data for model19 .. taken with
- 30oct17: no td tracking, 430lf ballast added to az arm
- 22,24,25,26oct17: no td tracking, pre 430lf
- 20sep17: hurricane maria
- 01aug17: cband
wide installed. cal values and optical alignment are
- 13jun15: lbw changed 1295 freq -> 1310 to get away
from punta salinas new freq, changed l175 -> 1150 to get away
(file .2). lbw low band was taken with circular rather
than linear mode.
- 21may14: alfa reinstalled.
- 28mar14- alfa down to lab. weight change may affect pointing
- 13jan13: updated pointing error offsets using jan12-dec12
- 27feb12-04mar12 model 18 taken with sbn. The pointing error
will be large since the tiedowns used a new position to remove
the residual platform tilt.
- nov10: cband Tsys increased to about 40k on 13nov10 and then
went back down around about 22nov10.
- Caused by the gortex horn cover coming off and moisture
getting in the mouth of the horn.
- 03jun10: updated
model with 1,2,3 az fits to data mar10-jun10
- 24mar10: repaired broken beam, za limits 3.5 to 12 deg za
- 02mar09: switched the pointing errors to be from 01jul08..
last mod to pointing model.
- 01jul08: xband model subtracted 4 asecs from za constant term
for bania. Did not modify any other recievers.
updated pointing model with calibration data using fits to
1az and 3 az terms.
- 13jun08: cooled 327 receiver installed
- dec07: this included the pointing model made with sband narrow
after the painting project completed. It will include the
pointing errors after the painting and before the new model was
- xxapr07: pointing off since painting program moving weights.
- 31mar07: painting project started moving weight on the
- 10dec06: updated pointing model 1az terms and az*sinza.
- 03/04dec06: x102 runs had offsets in model for testing.
sin(za) az: -45.9*sin(za). 1az term for za: -6.7cos(az) -
6.8sin(az) . These values were added to model. To goback to old
model, subtract these values.
- 16jun06: lbw polB Cal. found loose sma feed thru inside dewar.
Probably causing the drifts in polB Tsys.
- 10apr06: sbn cal remeasured 15mar06. installed 10apr06.
backdated to 14oct03. Need to recompute the sbn data to used
these new cal values. the old ones were probably a little high.
- 04dec05: lbw polB cal value jumps 20%. remeaured cals on
- 23nov05: lbw diodes replaced
- 30sep05: station clock off by 1 second midnite to
8:30am. lbw calibration data will have extra 15 asec pnterr.
- 18aug05: sbw cal cables changed. This changed the amount of
cal getting into the dewar. Need to remeasure the cals.
- 08aug05: new cband cals installed. backdated to
25may05 when new amps were installed. recomputed cband data from
may05 with new cal values.
- 22jul05: lbw diode 1 died.. used diode 2 for
cals until new diodes arrive.
- 21jun05: installed new lbw cals (had a bug in
diode1->polb... fixed 09aug05).
- 01jun05 : lbw: found bad polA connection to
probe in dewar. found cracke polB cal cable in dewar..
- 25may05: cband reinstalled with new amps, cal
values need to be updated.
- 13mar05 Included 430ch data in system
- 12mar05 oh filterbank installed for alfa.
- 29jan05 (check date) 430 filters to
remove tv station intermods installed polb ch and pola,b dome.
- 23jan05 recomputed alfa system performance
using the new cal values.
- 05jan05 added alfa data. Using old cal
values from the lab.
- 27dec04 cbh . model offsets az,za
incremented by (-.28,-3.82). data from x102 runs.
- 27oct04 the 327 data taken 27,28oct04 had a
hybrid in front of the receiver. It was used to create circular
polarizations for a vlbi observation.
- 09oct04 cband dewar brought down. found
short in polA cable inside dewar that was causing the tsys
values to jump around.
- 12sep04 lbw cals measured 06aug04, installed
12sep04, backdated to 04jun04.
- 04jul04 cband polA cal value jumps back
(after dewar warmed up and cooled down). note continues to jump
around until 08oct04 when dewar brought down and short found in
one of the cables in the dewar.
- 31may04 cband polA cal value jumps.
- 17may04 pointing model 15 installed after
verify runs completed. (sbn installed a few days before).
- 14may04 sbw diodes replaced, horn lowered
- 10may04 during this week calib scans were done
with some receivers with large pointing errors.
- 10may04 during week pointing model made sbn,
verification other sources with calib scans.
- 01may04 tertiary resurveyed into position
- 29apr04 horns resurveyed and moved into
"best" position. lbw,sbw,sbn not moved.
- 23apr04 lbw
polB cal value jumps.
- 21apr04 alfa installed, kevlar cables installed
- 15mar04 xb receiver back up. Was in lab for
cals,amps, loose cables in dewar. remeasured cal values.
- 29feb04 pointing offsets cbh updated using
calib runs feb04.
- 14feb04 horn survey, repositioned
horns, updated pointing model. Larges move was lbw.
- 13feb04 sbh horn moved dow .8 inches to
bring in focus. New cals installed
- 10feb04 cbh installed
- 26jan04 leaf springs on the dome changed to
stiffer ones (approx. date)
- 14jan04 610 receiver moved on the turret
- 07jan04 327,xband receivers moved on the
- 10oct03 sbw cals remeasured. backdated to
15aug03. still drifting.
- 02nov03 lbw gain curve from data 01mar03 thru sep03.
back dated to 01mar03.
- 14oct03 8am. az encoder jumped by 3 teeth
(about 16 asecs). fixed on 16oct03.
- 05oct03 xband
cals remeasured on antenna test rangei valid from 05oct03.
- 04sep03 I think the Cal values on xband have
changed after new horn installed (see
measurements).. They need to be remeasured.
- 03sep03 new xband horn installed.
- 21jul03 xband. updated pointing offsets from
- 21jul03 xband installed new cal values from
- 30may03 recomputed xband cals. uncorcal,0cal definitions were
wrong on measurement page.
- 18may03 loose cable cband cals. retightened cable and redid
- 14may03 remeasured sbh cal values. diodes have been drifting.
- 22apr03: cband new cal noise diodes installed.
- 10mar03: corrected feeds offsets installed 07mar03.
- 07mar03: model 14 installed with feeds offsets (but offsets
- xxfeb03: dome lifted, elevation rails shimmed, model 14
- 21nov02: lbw horn focus changed
- 30aug03: 610 antenna switched from fat dipole back to skinny
- 28aug02: cband horn focus changed
- 21aug02: lbn horn focus changed.
- 31jul02: 327 room temperature receiver installed.
- 02apr02: update xband turret position from theodolite survey.
- 23may01: 610 receiver switched to fat dipole.
- 06feb02: reshimmed xband horn position.
- xxsep01: major adjustments to reflector surface completed.
Misc, multi-freq: (top)
27feb18: Gain variations
before/after hurricane maria.
sysperf jun12-dec12. updated pointing offsets.
24mar12: gain vs freq and
ruze curves (data mar11 thru mar12)
10mar09: gain vs frequency and
the ruze curves show the surface errors.
14feb05: hpbw vs lambda (all
rcvrs) for data taken 2004
14feb05: hi frequency performance
after reflector adjustments.